Find the most general antiderivative of f(x) = (3 – 4x) -3 – (1 + x2)-1
Answer #1

Giver f(x) = (3-44) 3 (1+22). S f(x) dx = f((3- 4 x 5 3 (1+x2)) = {(3-425² de- f it ya da As 8x dx = xht nti arctan (*) +C (